OpenStudy (anonymous):

if the sum of the zeros of rx^2+2x+3k is equal to their product.find the value of k.

OpenStudy (anonymous):

-2/r=3k/r k=-2/3

OpenStudy (watchmath):

:)

OpenStudy (anonymous):

m+n=-b/a and mn=c/a

OpenStudy (anonymous):

here a=r.b=2.c=3k and m +n=mn

OpenStudy (anonymous):

how k=_2/3

OpenStudy (anonymous):

-2/r=3k/r 3k=-2 k=-2/3